OpenStudy (anonymous):

Will give a medal and fan... Mike was working on solving the exponential equation 37x = 12; however, he is not quite sure where to start. Using complete sentences, describe to Mike how to solve this equation and how solving would be different if the bases were equal.

OpenStudy (dumbcow):

If \[a^x = b\] then \[x = \log_{a} b\] If bases are equal \[a^x = a^n\] then you can just set exponents equal \[x = n\]

OpenStudy (anonymous):

So, 37x=12 Then, x=log37 (12) What do you mean if the bases are equal?

OpenStudy (dumbcow):

\[37^x = 37^4\] you dont have to use log just say x = 4

OpenStudy (anonymous):

Ohhhh okay! That makes so much more sense. Thank you!!! :D

OpenStudy (dumbcow):

yw :)